# Bias in Motion: Theoretical Insights into the Dynamics of Bias in SGD Training

### v1

@@ 4.1 Spurious correlations
From Eqs. 10-3.4, two relevant timescales for the problem are observed:
$\tau_{M}=\frac{1}{\eta(v+\Delta^{mix})},$ (13)
$\tau_{R}=1/\eta\Delta^{mix}.$ (14)
The shortest timescale, $\tau_{M}$, associated with $M$, indicates that the student first aligns with the spurious feature. By aligning with the shift vector, the student can predict most examples correctly, but not all. The effect of spurious correlations is transient; at $t\sim\tau_{R}$, the student starts disaligning from the spurious feature and aligns with the teacher vector, eventually achieving nearly perfect alignment. This is illustrated in Fig. 2, where the student initially picks up on the spurious correlation (green) and achieves almost perfect alignment with the shift vector during intermediate times before aligning nearly perfectly with the teacher (red).
@@ 4.2 Fairness
Zero shift. We first consider a simplified case where we assume that both clusters are centered at the origin $v=0$ as shown in Fig. 1c. We will later reintroduce the shift and analyse the transient dynamics it introduces as per the discussion in section 4.1. The zero shift case represents an extreme situation where not only is the classification not solvable if $\overline{\boldsymbol{w}}_{+}\neq\overline{\boldsymbol{w}}_{-}$, but it is also difficult to identify which cluster generated a given data point since the shift provides no information. This setting is particularly suited to analysing the effects of ‘group level’ features, such as group variance and relative representation, on the preference of the classifier.
@@
General case. We now consider the general case shown in Fig. 1d, where the shift is non zero and all three timescales identified so far play a role.
@@ 5 Ablations using numerical simulations
Figure 5: Numerical simulations on MNIST. The figure shows the average (solid lines) and standard deviation (shaded area) of 100 simulations run in this framework. In particular the upper plots show the test loss and lower plots the test accuracy for subpopulation $+$ (blue) and -- (red). Panel (a) an example of crossing phenomenon obtained by imposing $\sqrt{\Delta}_{+}=1$, $\sqrt{\Delta}_{-}=0.2$, and $\rho=0.1$. Panel (b) shows the double crossing, obtained by introducing an additional timescale to the previous case by tuning label imbalance.
Rotated MNIST. We train a 2-layer neural network with 200 hidden units, ReLU activation, and sigmoidal readout activation, using online SGD on MSE loss in a MNIST classification task. Data are centered and the variance is set to 1 following standard pre-processing practices. We consider a variation of the MNIST dataset that mimics the TM model, allowing us to verify our theory when network and the data structure are mismatched. Digits 0 to 4 and 5 to 9 are grouped to form the two subpopulations. With probability $p_{+}$ and $p_{-}$, digits of both subpopulations are rotated with a subpopulation-specific angle—i.e. Fig. 5a uses angles of rotation $\theta_{-}=45^{o}$ and $\theta_{-}=-90^{o}$. The goal of the classifier is to detect rotations.
@@
Finally, we consider the general fairness case. By creating label imbalance, i.e. setting $p_{+}=0.3$ and $p_{-}=0.7$, we observe an additional phase of bias evolution, wherein the classifier prefers dense regions with consistent labels. This advantages subpopulation -- and indeed it is what we see in Fig. 5b. The result of the simulations matches the theory displaying a double crossing phenomenon.
(a) Group brightness CIFAR10
(b) Crossing in CIFAR10
Figure 6: Numerical simulations on CIFAR10. The figure shows experiments of a 2L neural network on CIFAR10 where classes were grouped together to form the subpopulations. The plots show the average performance—measure by loss or accuracy—achieved over 100 simulations (for Panel (a)) and 10 simulations (for Panel (b), respectively) using the shaded area to quantify the standard deviation. Panel (a) shows the result at the end of training changing relative representation $\rho$, while Panel (b) shows the training trajectories in a particular instance, see text for more details.
CIFAR10. We consider the same architecture and pre-processing described for MNIST on a CIFAR10 classification task. We select 8 classes and assign 4 of them to the positive group and 4 to the negative group. Inside each group, 2 classes are labelled as negative and 2 as positive. This simulation framework is similar to the one considered by [5] where the authors used sub-populations with only 2 classes each.
@@
When labels are balanced, our theory predicts that the classifier is initially attracted by the larger $\Delta$ and eventually—if the relative representation of the group with smaller $\Delta$ is large enough—it switches and favours the other group. This effect is indeed verified in the CIFAR10 experiments. Starting from the partitioning in Fig. 6a (bottom) with $\rho=0.8$, the dynamics is initially attracted by group – before advantaging the other group, giving rise to a crossing as shown in Fig. 6b.
Additional numerical expirements. In Appendix F, we provide additional experiments within the TM model and the CelebA dataset, exploring different architectures and losses. Even under these new conditions, we observe that bias presents different timescales and shows crossing behaviors.
@@ Acknowledgements
We thank Andrew Saxe and Simon Lacoste-Julien for numerous discussions and insightful comments. This work was supported by the Sainsbury Wellcome Centre Core Grant from Wellcome (219627/Z/19/Z) and the Gatsby Charitable Foundation (GAT3755).
@@ F.2 Simulations on Synthetic Data and Deeper Networks
Figure 10: Simulations on Synthetic Data and Deeper Networks We observe the ‘double-crossing’ phenomena in not only the loss curves, but also the error curves for the positive sub-population (blue) and the negative sub-population (red). The shaded areas quantify the standard deviation obtained across 10 seeds. The data distribution parameters are $d=100,v=4,\rho=0.75,\Delta_{+}=0.1,\Delta_{-}=1,T_{\pm}=0.9,\eta=0.01,\alpha_{+}=0.473,\alpha_{-}=-0.200$
In this section we test the validity of the prediction of our model in more realistic settings. Specifically, assuming the same data distribution, we now train a multilayer perceptron (MLP) having one hidden layer of 200 units. We use ReLU activation and a sigmoid activation on the output. We train using online stochastic gradient descent and use binary cross entropy as our loss function. We sample training and test data from the data distribution and use the test data to obtain estimates of the loss as well as error rates (percentage of test examples misclassified).
For the general fairness case (sec.  4.2), we observe the three phase behaviour predicted by our model. The positive sub-population is initially advantaged more since it exhibits stronger spurious correlation. Then, the negative sub-population is advantaged since it has a higher variance. Finally, as per Eq. 16, the positive-sub-population is advantaged once more since it has sufficiently high representation. We not only observe the ‘double-crossing’ phenomena in the losses, but also in the test errors demonstrating the robustness of our model beyond the linearity and MSE loss assumptions.

### v2

@@ 4.1 Spurious correlations
From Eqs. 10-3.4, two relevant timescales for the problem are observed:my very long text $\tau_{M}=\frac{1}{\eta(v+\Delta^{mix})},$ (13) $\tau_{R}=1/\eta\Delta^{mix}.$ (14) The shortest timescale, $\tau_{M}$, associated with $M$, indicates that the student first aligns with the spurious feature. By aligning with the shift vector, the student can predict most examples correctly, but not all. The effect of spurious correlations is transient; at $t\sim\tau_{R}$, the student starts disaligning from the spurious feature and aligns with the teacher vector, eventually achieving nearly perfect alignment. This is illustrated in Fig. 2, where the student initially picks up on the spurious correlation (green) and achieves almost perfect alignment with the shift vector during intermediate times before aligning nearly perfectly with the teacher (red).
@@ 4.2 Fairness
4.2.1 Zero shift
We first consider a simplified case where we assume that both clusters are centered at the origin $v=0$ as shown in Fig. 1c. We will later reintroduce the shift and analyse the transient dynamics it introduces as per the discussion in section 4.1. The zero shift case represents an extreme situation where not only is the classification not solvable if $\overline{\boldsymbol{w}}_{+}\neq\overline{\boldsymbol{w}}_{-}$, but it is also difficult to identify which cluster generated a given data point since the shift provides no information. This setting is particularly suited to analysing the effects of ‘group level’ features, such as group variance and relative representation, on the preference of the classifier.
@@ 4.2.2 General case
We now consider the general case shown in Fig. 1d, where the shift is non zero and all three timescales identified so far play a role.
@@ 5.1 Rotated MNIST
(c) Effect of $\Delta_{-}$ on the initial learning phase
Figure 5: Numerical simulations on MNIST. The figure shows the average (solid lines) and standard deviation (shaded area) of 100 simulations run in this framework. In particular the upper plots show the test loss and lower plots the test accuracy for subpopulation $+$ (blue) and -- (red). Panel (a) an example of crossing phenomenon obtained by imposing $\sqrt{\Delta}_{+}=1$, $\sqrt{\Delta}_{-}=0.2$, and $\rho=0.1$. Panel (b) shows the double crossing, obtained by introducing an additional timescale to the previous case by tuning label imbalance. Panel (c) explore the effect of changing $\Delta_{-}$ while keeping a constant $\Delta_{+}=1$.
We train a 2-layer neural network with 200 hidden units, ReLU activation, and sigmoidal readout activation, using online SGD on MSE loss in a MNIST classification task. Data are centered and the variance is set to 1 following standard pre-processing practices. We consider a variation of the MNIST dataset that mimics the TM model, allowing us to verify our theory when network and the data structure are mismatched. Digits 0 to 4 and 5 to 9 are grouped to form the two subpopulations. With probability $p_{+}$ and $p_{-}$, digits of both subpopulations are rotated with a subpopulation-specific angle—i.e. Fig. 5a uses angles of rotation $\theta_{-}=45^{o}$ and $\theta_{-}=-90^{o}$. The goal of the classifier is to detect rotations.
@@
In Fig. 5b, we extend the MNIST experiment by varying the average image brightness of subpopulation --, which reflects the group variance in our theory. The results show that greater brightness leads to faster learning in the second phase and an increasing asymptotic preference, consistent with our predictions.
Finally, we consider the general fairness case. By creating label imbalance, i.e. setting $p_{+}=0.3$ and $p_{-}=0.7$, we observe an additional phase of bias evolution, wherein the classifier prefers dense regions with consistent labels. This advantages subpopulation -- and indeed it is what we see in Fig. 5c. The result of the simulations matches the theory displaying a double crossing phenomenon.
5.2 CIFAR10
(a) Group brightness
(b) Crossing and $\rho$ dependence
Figure 6: Numerical simulations on CIFAR10. The figure shows experiments of a 2L neural network on CIFAR10 where classes were grouped together to form the subpopulations. The plots show the average performance—measure by loss or accuracy—achieved over 100 simulations (for Panel (a)) and 10 simulations (for Panel (b), respectively) using the shaded area to quantify the standard deviation. Panel (a) shows the result at the end of training changing relative representation $\rho$, while Panel (b) shows the training trajectories as $\rho$ changes as indicated in the colour-bar, see text for more details.
We consider the same architecture and pre-processing described for MNIST on a CIFAR10 classification task. We select 8 classes and assign 4 of them to the positive group and 4 to the negative group. Inside each group, 2 classes are labelled as negative and 2 as positive. This simulation framework is similar to the one considered by [5] where the authors used sub-populations with only 2 classes each.
@@ 5.2 CIFAR10
When labels are balanced, our theory predicts that the classifier is initially attracted by the larger $\Delta$ and eventually—if the relative representation of the group with smaller $\Delta$ is large enough—it switches and favours the other group. This effect is indeed verified in the CIFAR10 experiments. Starting from the partitioning in Fig. 6a (bottom) with $\rho=0.8$, the dynamics is initially attracted by group – before advantaging the other group, giving rise to a crossing that can be observed–among other things–in Fig. 6b.
Fig.6b highlights another prediction of our theory concerning the timescale when $\rho$ becomes relevant. Our theory predicts that $\rho$ should become relevant only in the final stages of the dynamics and indeed the curves are almost perfectly overlapping until for most of training. As already noticed, $\rho$ sufficiently large gives rise to a crossing behaviour as indicated demonstrated in the synthetic model.
5.3 Additional numerical experiments
In Appendix F, we provide additional experiments within the TM model and the CelebA dataset, exploring different architectures and losses. Even under these new conditions, we observe that bias presents different timescales and shows crossing behaviours.
@@ Acknowledgements
We thank Andrew Saxe and Simon Lacoste-Julien for numerous discussions and insightful comments. Anchit Jain is grateful to the Gatsby Computational Neuroscience Unit for the research internship opportunity. This work was supported by the Sainsbury Wellcome Centre Core Grant from Wellcome (219627/Z/19/Z) and the Gatsby Charitable Foundation (GAT3755).
@@ F.2 Simulations on Synthetic Data and Deeper Networks
Figure 10: Synthetic Data Simulation with alternate Training Protocols We observe the ‘double-crossing’ phenomena in not only the loss curves, but also the error curves for the positive sub-population (blue) and the negative sub-population (red) (left). The shaded areas quantify the standard deviation obtained across 10 seeds. We observe similar behavior when using Adam (middle) and weight decay (right). The data distribution parameters are $d=100,v=4,\rho=0.7,\Delta_{+}=0.1,\Delta_{-}=1,T_{\pm}=0.9,\eta=0.01,\alpha_{+}=0.471,\alpha_{-}=-0.188$
In this section we test the validity of the prediction of our model across various networks and training protocols. We consider a data distribution with parameters as detailed in the caption of Fig. 10. We train a 2-layer MLP with ReLU activation in the hidden layer(s) and sigmoidal activation at the output with 128 neurons in the hidden layer with online SGD using BCE loss. We refer to this as the ‘standard configuration’ for further comparisons. We sample training and test data from the data distribution and use the test data to obtain estimates of the loss as well as error rates (percentage of test examples misclassified). We visualize these in Fig. 10 on the left.
We observe the three phase behaviour predicted by our model. The positive sub-population is initially advantaged more since it exhibits stronger spurious correlation. Then, the negative sub-population is advantaged since it has a higher variance. Finally, as per Eq. 16, the positive-sub-population is advantaged once more since it has sufficiently high representation. We not only observe the ‘double-crossing’ phenomena in the losses, but also in the test errors demonstrating the robustness of our model beyond the linearity and MSE loss assumptions.
We also observe similar behavior using Adam optimization in Fig. 10 (center). We note that we had to use a smaller learning rate of 0.001 to keep training stable since Adam leads to faster optimization manifesting in a higher effective learning rate than SGD.
When we train using a weight decay penalty of 0.1 (Fig. 10 right), we observe that asymptotically the higher variance cluster is now preferred. This behaviour can be explained using the theoretical model. As discussed in Appendix  E, smaller student magnitude advantages the higher variance group and weight decay encourages smaller student weights leading to an asymptotic preference for the negative sub-population.
Figure 11: Ablations with Deeper Networks We observe the ‘double-crossing’ phenomena across deeper networks as well.
Figure 12: Ablations with Wider Networks We observe the ‘double-crossing’ phenomena across wider networks as well.
Figure 13: Ablations across Learning Rates Larger learning rates can lead to instability (left). If training is stable however, we observe the ‘crossing’ phenomena as usual, just at different time scales due to different speeds of training.
We now train deeper networks with 2 to 5 layers and visualise results in Fig. 11. The ‘double-crossing’ phenomena persists across deeper networks. We also train wider networks with 2, 16, 128 and 1024 units in the hidden layer and visualise results in Fig. 12. The ‘double-crossing’ phenomena persists across wider networks.
Finally, we vary the learning rate as well from our standard configuration and note that when the learning rate is too high, the training is unstable for the higher variance cluster (Fig. 13 left). For stable training, however, we observe the ‘double-crossing’ phenomena as usual, just at longer timescales for slower learning rates as predicted by our theory.